a car traveling at a constant speed travel 90km in 50 minutes. how far, traveling at the same constant speed, will the car travel in 2 hour 25 minutes?

Answer by Alan3354(69443)

a car traveling at a constant speed travel 90km in 50 minutes. how far, traveling at the same constant speed, will the car travel in 2 hour 25 minutes?
----------
Find its speed:
r = d/t = 90/50 = 1.8 km/minute
------------
2:25 = 145 minutes
d = r*t